OpenStudy (anonymous):

Find the value of x and y when matrix[3x,2y,6y] = [x+5,0,10]. Thanks!

OpenStudy (ghazi):

equate the corresponding parts... 3x=x+5 2y=0 and 6z=10

OpenStudy (ghazi):

check out your matrix again

OpenStudy (anonymous):

Oh, there's a typo. it's [3y,2x,6y]. Sorry.

OpenStudy (ghazi):

3y=x+5 2x=0 6y=10..therefore x=0 from second equation

OpenStudy (anonymous):

Thanks.

OpenStudy (ghazi):

y=5/3